Solve the following nonlinear system for x and y. Sketch a graph, showing the intersection points.
x + y = 7
x^2 + y^2 = 25



Answer by pushpaharan(47)

You need graphing calculator or you need to sketch it to solve.
x + y = 7 = linear equation
x^2 + y^2 = 25 - circle with center(0,0) radius 5
or you can solve algebrically.
y= 7-x
x^2+(7-x)^2=25
x^2+ 49-14x+x^2 =25
2x^2-14x+24=0
x^2 -7x +12 = 0
(x-4)(x-3)=0
therefore x=3 or x=4
y=4 y=3